OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 14x + 18y = -54 -4x – 14y = 42

OpenStudy (anonymous):

multiply second equation with 7 hen add both -14x-98y=294 now add this to first equation -80y=240 y=240/(-80) y=-3 put this in any of the above equations to get x

OpenStudy (radar):

Multiplying the second equation by 7 gives me -28x -98y=294 Not -14x-98y = 294

OpenStudy (radar):

A good start, but will require multiplying the first equation by 2. This will give you 28x+36y=-108

OpenStudy (radar):

Now you can eliminate x by combining: -28x-98y=294 28x+36y=-108 -62y=186 y=-3 same as sami-21

OpenStudy (radar):

Solve for x by substituting the -3 for y in one of the original equations. 14x+18(-3)=-54 14x=54-54=0 x=0

OpenStudy (radar):

Since x=0, The error in a previous post did not affect the value for y.

OpenStudy (anonymous):

what?? your confusing me now

OpenStudy (radar):

Read all the post and your confusion will disappear.

OpenStudy (radar):

Pay particular to the very first post by sam-21, do you agree with the results that were obtained by multiplying the second given equation by 7, did you get that?

OpenStudy (radar):

Do you see how he eliminated x that way (without multiplying the first equation by 2)?

OpenStudy (anonymous):

yes

OpenStudy (radar):

What do you get when you multiply the second equation by 7 (-4x-14y=42), I get -28x-98y=294 what do you get?

OpenStudy (radar):

What did sami-21 get? See the difference..??